Solve the system of equations
5x + 2y = 8
12x + 6y = 20

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

Solve the system of equations
5x + 2y = 8
12x + 6y = 20
====================
Multiply the 1st eqn by 3, then subtract

Answer by jim_thompson5910(35256)

Hint: Multiply both sides of equation 1 by -3 to get -15x-6y = -24

Then add this to equation 2

-15x-6y = -24
12x + 6y = 20
--------------
-3x + 0y = -4


so -3x = -4

solve for x, then use this to find y.